OpenStudy (anonymous):

x sin y + cos 2y = cos y

OpenStudy (lgbasallote):

nice equation....

OpenStudy (anonymous):

I'm sorry I left out some stuff there I am trying to find dy/dx using implicit differentiation

OpenStudy (dape):

Differentiate both sides with respect to x, then solve for dy/dx.

OpenStudy (anonymous):

How would I go about that something like cos x y+cos 2y +sin x y + -sin y=-sin y?

OpenStudy (anonymous):

I know i'm leaving something out but not sure what

OpenStudy (dape):

Not sure what you are asking, could you be more specific?

OpenStudy (dape):

So what you want to do is to solve the following equation for dy/dx: \[ \frac{d}{dx}(x\sin{y}+\cos{2y})=\frac{d}{dx}\cos{y} \]

OpenStudy (anonymous):

I am trying to find dy/dx using implicit differentiation for the equation x sin y+ cos 2y = cos y was that more helpful?

OpenStudy (dape):

So for the left side you have: \[ \frac{d}{dx}(x\sin y+\cos 2y)=\frac{d}{dx}x\sin y+\frac{d}{dx}\cos 2y \\ = \sin y+x\cos(y)\frac{dy}{dx}-\sin(2y)*2\frac{dy}{dx}\]

OpenStudy (dape):

Now just do the right side and solve for dy/dx like it's a normal equation.

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (dape):

No problem! =)